What is the area of the circumscribing circle if the side of a right angle triangle are 3,4, and 5 meter?
1
Expert's answer
2020-09-22T15:59:44-0400

Explanations & Calculations


  • Circumscribed circle is a one drawn though all the vertices of a polygon.
  • The aim is to find the radius of the circle so that it's possible to calculate its area (A).
  • One should know that, it's a special occasion when it comes to a right triangle for which the circumscribing circle to be drawn and the hypotenuse of the triangle becomes the diameter of the circle hence the radius is the half of it.
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small A &= \small \pi r^2\\
&= \small \pi \times \Big(\frac{5}{2}\Big)^2\\
&= \small \bold{19.635\,m^2}
\end{aligned}"